A 0.435 kg blue bead slides on a curved frictionle
Answer #1

mgh = 1/mv^2

v = sqrt(2gh)

v = sqrt(2*9.8*1.5) = 5.42 m/s

the velocity v2' of the ball after collision is found by
v2' = 2(m1v1 + m2v2)/(m1+m2) - v2

v2 = initial speed of the ball= 0,
v1 = initial speed of the bead

v2' = 2(0.435*5.42)/1.04 = 4.53 m/s

now

1/2mv2'^2 = mgh'

h' = 4.53^2/2*9.8

h' = 1.047 m